Use the formula C= 5/9 (F-32 degrees) to find the Celsius temperature that is equivalent to the 59 degrees Fahrenheit.
Answer by checkley71(8403)

C=5/9(F-32)
C=5/9(59-32)
C=5/9*27
C=5*3
C=15 ANSWER.